# Factor Augmented Tensor-on-Tensor Neural Networks

### v1

@@ A.1 Notation
which equals to the sine of the largest principle angle between the column spaces of $U$ and $\widehat{U}$. For any two matrices $A\in\mathbb{R}^{m_{1}\times r_{1}},B\in\mathbb{R}^{m_{2}\times r_{2}}$, denote the Kronecker product $\odot$ as $A\odot B\in\mathbb{R}^{m_{1}m_{2}\times r_{1}r_{2}}$. For any two tensors ${\cal A}\in\mathbb{R}^{m_{1}\times m_{2}\times\cdots\times m_{K}},{\cal B}\in\mathbb{R}^{r_{1}\times r_{2}\times\cdots\times r_{N}}$, denote the tensor product $\otimes$ as ${\cal A}\otimes{\cal B}\in\mathbb{R}^{m_{1}\times\cdots\times m_{K}\times r_{1}\times\cdots\times r_{N}}$, such that
@@
A.2 Preliminary of matrix and tensor algebra
@@
Lemma A.1 (Weyl’s inequality).
@@
$|\sigma_{i}(A)-\sigma_{i}(B)|\leq\|A-B\|$ (S.2)
Lemma A.2 (Davis–Kahan $\sin\Theta$ theorem).
@@
A.3 Technical lemmas
Lemma A.3 ($\epsilon$-covering of matrix norms, Han et al., (2020)).
@@ A.3 Technical lemmas
$\max_{M\in\mathbb{R}^{d\times d^{\prime}},\|M\|_{2}\leq 1,\text{rank}(M)\leq r}\ \min_{j\leq N_{dr,\epsilon/2},j^{\prime}\leq N_{d^{\prime}r,\epsilon/2}}\|M-U_{j}V_{j^{\prime}}^{\top}\|_{2}\leq\epsilon.$
@@
$\sup_{M,\widetilde{M}\in\mathbb{R}^{d\times d^{\prime}},\|M-\widetilde{M}\|_{2}\leq\epsilon\atop{\|M\|_{2}\vee\|\widetilde{M}\|_{2}\leq 1\atop\text{rank}(M)\vee\text{rank}(\widetilde{M})\leq r}}\frac{\|f(M-\widetilde{M})\|_{*}}{\epsilon 2^{I_{r<d\wedge d^{\prime}}}}\leq\sup_{\|M\|_{2}\leq 1\atop\text{rank}(M)\leq r}\|f(M)\|_{*}\leq 2\max_{1\leq j\leq N_{dr,1/8}\atop 1\leq j^{\prime}\leq N_{d^{\prime}r,1/8}}\|f(U_{j}V_{j^{\prime}}^{\top})\|_{*}.$ (S.5)
(iii) Given $\epsilon>0$, there exist $U_{j,k}\in\mathbb{R}^{d_{k}\times r_{k}}$ and $V_{j^{\prime},k}\in\mathbb{R}^{d^{\prime}_{k}\times r_{k}}$ with $\|U_{j,k}\|_{2}\vee\|V_{j^{\prime},k}\|_{2}\leq 1$ such that
$\max_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\|M_{k}\|_{2}\leq 1\atop\text{rank}(M_{k})\leq r_{k},\forall k\leq K}\ \min_{j_{k}\leq N_{d_{k}r_{k},\epsilon/2}\atop j^{\prime}_{k}\leq N_{d_{k}^{\prime}r_{k},\epsilon/2},\forall k\leq K}\Big{\|}\odot_{k=2}^{K}M_{k}-\odot_{k=2}^{K}(U_{j_{k},k}V_{j_{k}^{\prime},k}^{\top})\Big{\|}_{\rm op}\leq\epsilon(K-1).$
@@
$\sup_{M_{k},\widetilde{M}_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\|M_{k}-\widetilde{M}_{k}\|_{2}\leq\epsilon\atop{\text{rank}(M_{k})\vee\text{rank}(\widetilde{M}_{k})\leq r_{k}\atop\|M_{k}\|_{2}\vee\|\widetilde{M}_{k}\|_{2}\leq 1\ \forall k\leq K}}\frac{\|f(\odot_{k=2}^{K}M_{k}-\odot_{k=2}^{K}\widetilde{M}_{k})\|_{*}}{\epsilon(2K-2)}\leq\sup_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}}\atop{\text{rank}(M_{k})\leq r_{k}\atop\|M_{k}\|_{2}\leq 1,\forall k}}\Big{\|}f\big{(}\odot_{k=2}^{K}M_{k}\big{)}\Big{\|}_{*}$ (S.6)
@@
$\sup_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\|M_{k}\|_{2}\leq 1\atop\text{rank}(M_{k})\leq r_{k}\ \forall k\leq K}\Big{\|}f\big{(}\odot_{k=2}^{K}M_{k}\big{)}\Big{\|}_{*}\leq 2\max_{1\leq j_{k}\leq N_{d_{k}r_{k},1/(8K-8)}\atop 1\leq j_{k}^{\prime}\leq N_{d_{k}^{\prime}r_{k},1/(8K-8)}}\Big{\|}f\big{(}\odot_{k=2}^{K}U_{j_{k},k}V_{j_{k}^{\prime},k}^{\top}\big{)}\Big{\|}_{*}.$ (S.7)
Lemma A.4 (Slepian’s inequality, Vershynin, (2018)).
@@
$\mathbb{E}\sup_{t\in T}X_{t}\leq\mathbb{E}\sup_{t\in T}Y_{t}$ (S.10)
Lemma A.5 (Sudakov-Fernique inequality, Vershynin, (2018)).
@@
Lemma A.6 (Gaussian Concentration of Lipschitz functions, Vershynin, (2018)).
@@
Lemma A.7 (Sigular values of Gaussian random matrices).
Let $X$ be an $n\times p$ matrix with i.i.d. $N(0,1)$ entries. For any $x>0$,
$\mathbb{P}\left(\left\|X^{\top}X-nI_{p}\right\|_{2}\geq p+2\sqrt{pn}+2x(\sqrt{p}+\sqrt{n})+x^{2}\right)\leq 2e^{-x^{2}/2}.$
@@
It can be easily derived by Corollary 5.35 in Vershynin, (2010). ∎
A.4 The Determination of Rank
Here, the estimators are constructed with given ranks $r_{1},\ldots,r_{K}$. Determining the number of factors in a data-driven manner has been a significant research topic in the factor model literature. Recently, Han et al., (2022) established a class of rank determination approaches for tensor factor models, based on both the information criterion and the eigen-ratio criterion. These procedures can be adopted for our purposes.
A.5 Tucker Decomposition and iterative TIPUP (iTIPUP) Algorithm
The Higher-Order Orthogonal Iteration (HOOI) algorithm De Lathauwer et al., (2000) is one of the most popular methods for computing the Tucker decomposition, which relies on the orthogonal projection and Singular Value Decomposition (SVD) to update the estimation of loading matrices for each tensor mode in every iteration.

### v2

@@ Appendix A Appendix A: Tensor Factor Models
which equals to the sine of the largest principal angle between the column spaces of $U$ and $\widehat{U}$. For any two matrices $A\in\mathbb{R}^{m_{1}\times r_{1}},B\in\mathbb{R}^{m_{2}\times r_{2}}$, denote the Kronecker product $\odot$ as $A\odot B\in\mathbb{R}^{m_{1}m_{2}\times r_{1}r_{2}}$. For any two tensors ${\cal A}\in\mathbb{R}^{m_{1}\times m_{2}\times\cdots\times m_{K}},{\cal B}\in\mathbb{R}^{r_{1}\times r_{2}\times\cdots\times r_{N}}$, denote the tensor product $\otimes$ as ${\cal A}\otimes{\cal B}\in\mathbb{R}^{m_{1}\times\cdots\times m_{K}\times r_{1}\times\cdots\times r_{N}}$, such that
@@
For vectors $a_{1},\ldots,a_{K}$ of length $r_{1},\ldots,r_{K}$, the outer product
Let $M_{1}\in\mathbb{R}^{r_{1}\times R},\ldots M_{K}\in\mathbb{R}^{r_{K}\times R}$ be $K$ matrices with the same column dimension (denoted by $R$), the notation $\llbracket M_{1},\cdots,M_{K}\rrbracket$ represents
where $m_{kr}$ is the $r$th column of $M_{k}$, $k=1,\ldots,K$, $r=1,\ldots,R$.
For two tensor $\mathcal{A}:d_{1}\times\cdots\times d_{K}\times r_{1}\times\cdots\times r_{K}$ and $\mathcal{B}:r_{1}\times\cdots\times r_{K}\times p_{1}\times\cdots\times p_{q}$, the notation $\langle{\cal A},{\cal B}\rangle_{L}$ denote the contracted tensor product (Lock 2018)
[…]
Preliminary of matrix and tensor algebra
@@
Lemma 1 (Weyl’s inequality).
@@
$|\sigma_{i}(A)-\sigma_{i}(B)|\leq\|A-B\|.$ (S.4)
Lemma 2 (Davis–Kahan $\sin\Theta$ theorem).
@@
Technical lemmas
Lemma 3 ($\epsilon$-covering of matrix norms, Han et al. (2024a)).
@@
$\max_{M\in\mathbb{R}^{d\times d^{\prime}},\|M\|_{2}\leq 1,\text{rank}(M)\leq r}$
$\min_{j\leq N_{dr,\epsilon/2},j^{\prime}\leq N_{d^{\prime}r,\epsilon/2}}\|M-U_{j}V_{j^{\prime}}^{\top}\|_{2}\leq\epsilon.$
@@
$&\sup_{M,\widetilde{M}\in\mathbb{R}^{d\times d^{\prime}},\|M-\widetilde{M}\|_{2}\leq\epsilon\atop{\|M\|_{2}\vee\|\widetilde{M}\|_{2}\leq 1\atop\text{rank}(M)\vee\text{rank}(\widetilde{M})\leq r}}\frac{\|f(M-\widetilde{M})\|_{*}}{\epsilon 2^{I_{r<d\wedge d^{\prime}}}}\\ \leq&\sup_{\|M\|_{2}\leq 1\atop\text{rank}(M)\leq r}\|f(M)\|_{*}\leq 2\max_{1\leq j\leq N_{dr,1/8}\atop 1\leq j^{\prime}\leq N_{d^{\prime}r,1/8}}\|f(U_{j}V_{j^{\prime}}^{\top})\|_{*}.\end{aligned}$ (S.7)
(iii) Given $\epsilon>0$, there exist $U_{j,k}\in\mathbb{R}^{d_{k}\times r_{k}}$ and $V_{j^{\prime},k}\in\mathbb{R}^{d^{\prime}_{k}\times r_{k}}$ with $\|U_{j,k}\|_{2}\vee\|V_{j^{\prime},k}\|_{2}\leq 1$ such that $\forall k\leq K$
$\max_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\atop\|M_{k}\|_{2}\leq 1,\text{rank}(M_{k})\leq r_{k}}\min_{j_{k}\leq N_{d_{k}r_{k},\epsilon/2}\atop j^{\prime}_{k}\leq N_{d_{k}^{\prime}r_{k},\epsilon/2}}\Big{\|}\odot_{k=2}^{K}M_{k}-\odot_{k=2}^{K}(U_{j_{k},k}V_{j_{k}^{\prime},k}^{\top})\Big{\|}_{\rm op}$
$\leq\epsilon(K-1).$
@@
$&\sup_{M_{k},\widetilde{M}_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\|M_{k}-\widetilde{M}_{k}\|_{2}\leq\epsilon\atop{\text{rank}(M_{k})\vee\text{rank}(\widetilde{M}_{k})\leq r_{k}\atop\|M_{k}\|_{2}\vee\|\widetilde{M}_{k}\|_{2}\leq 1\ \forall k\leq K}}\frac{\|f(\odot_{k=2}^{K}M_{k}-\odot_{k=2}^{K}\widetilde{M}_{k})\|_{*}}{\epsilon(2K-2)}\\ &\leq\sup_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}}\atop{\text{rank}(M_{k})\leq r_{k}\atop\|M_{k}\|_{2}\leq 1,\forall k}}\Big{\|}f\big{(}\odot_{k=2}^{K}M_{k}\big{)}\Big{\|}_{*}\end{aligned}$ (S.8)
@@
$&\sup_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\|M_{k}\|_{2}\leq 1\atop\text{rank}(M_{k})\leq r_{k}\ \forall k\leq K}\Big{\|}f\big{(}\odot_{k=2}^{K}M_{k}\big{)}\Big{\|}_{*}\\ \leq&2\max_{1\leq j_{k}\leq N_{d_{k}r_{k},1/(8K-8)}\atop 1\leq j_{k}^{\prime}\leq N_{d_{k}^{\prime}r_{k},1/(8K-8)}}\Big{\|}f\big{(}\odot_{k=2}^{K}U_{j_{k},k}V_{j_{k}^{\prime},k}^{\top}\big{)}\Big{\|}_{*}.\end{aligned}$ (S.9)
Lemma 4 (Slepian’s inequality, Vershynin (2018)).
@@
$\mathbb{E}\sup_{t\in T}X_{t}\leq\mathbb{E}\sup_{t\in T}Y_{t}.$ (S.12)
Lemma 5 (Sudakov-Fernique inequality, Vershynin (2018)).
@@
Lemma 6 (Gaussian Concentration of Lipschitz functions, Vershynin (2018)).
@@
Lemma 7 (Sigular values of Gaussian random matrices).
Let $X$ be an $n\times p$ matrix with i.i.d. $N(0,1)$ entries. For any $x>0$, $\mathbb{P}\left(\left\|X^{\top}X-nI_{p}\right\|_{2}\geq p+2\sqrt{pn}+2x(\sqrt{p}+\sqrt{n})+x^{2}\right)\leq 2e^{-x^{2}/2}.$
@@
It can be easily derived by Corollary 5.35 in Vershynin (2010). ∎
The Determination of Rank
Here, the estimators are constructed with given ranks $r_{1},\ldots,r_{K}$. Determining the number of factors in a data-driven manner has been a significant research topic in the factor model literature. Recently, (Han, Chen, and Zhang 2022) established a class of rank determination approaches for tensor factor models, based on both the information criterion and the eigen-ratio criterion. These procedures can be adopted for our purposes.
Tucker Decomposition and iterative TIPUP (iTIPUP) Algorithm
The Higher-Order Orthogonal Iteration (HOOI) algorithm (De Lathauwer, De Moor, and Vandewalle 2000) is one of the most popular methods for computing the Tucker decomposition, which relies on the orthogonal projection and Singular Value Decomposition (SVD) to update the estimation of loading matrices for each tensor mode in every iteration.
@@
Proof of Proposition 1
Proof.
Let $L_{k}^{(m)}$ be the loss for $\hat{A}_{k}^{(m)}$ at the $m$-th iteration,
We outline the proof as follows.
(i) Initialization.
Define
The “noiseles” version of TIPUP1 is
which is of rank $r_{1}$. Then, by Davis-Kahan sin $\Theta$ theorem (Lemma 2),
To bound the numerator on the right-hand side of (S.15), we write
${\rm TIPUP}_{1}-\Theta_{1}$
$=$ $\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}{\cal E}_{t}^{\top}+\frac{1}{n}\sum_{t=1}^{n}{\cal E}_{t}A_{2}{\cal F}_{t}^{\top}A_{1}^{\top}+\frac{1}{n}\sum_{t=1}^{n}{\cal E}_{t}{\cal E}_{t}^{\top}$
::=:= $\Delta_{1}+\Delta_{2}+\Delta_{3}.$
For $\Delta_{3}$, by Lemma 7,
$\mathbb{P}\left(\left\|\sum_{t=1}^{n}\mathcal{E}_{t}\mathcal{E}_{t}^{\top}-n\mathbb{E}\mathcal{E}_{t}\mathcal{E}_{t}^{\top}\right\|_{2}\right.$
$\qquad\geq\left.d_{1}+2\sqrt{d_{1}d_{2}n}+2x(\sqrt{d_{1}}+\sqrt{d_{2}n})+x^{2}\right)$
$\leq$ $2e^{-x^{2}/2}.$
By setting $x\asymp\sqrt{d_{1}}$, with probability at least $1-e^{-d_{1}}$,
Thus for $u_{i}$ and $v_{i}$ with $\|u_{i}\|_{2}=\|v_{i}\|_{2}=1$, $i=1,2$,
$n\|\Theta_{1}\|_{2}^{-1}\overline{\mathbb{E}}\left((u_{1}^{\top}\Delta_{1}v_{1}-u_{2}^{\top}\Delta_{1}v_{2})^{2}\right)$
$\leq\left(\|u_{1}-u_{2}\|_{2}\|v_{1}\|_{2}+\|u_{2}\|_{2}\|v_{1}-v_{2}\|_{2}\right)^{2}$
$\leq 2\left(\|u_{1}-u_{2}\|_{2}^{2}+\|v_{1}-v_{2}\|_{2}^{2}\right)$
where $\xi$ and $\zeta$ are i.i.d $N(0,I_{d_{1}})$ vectors. As $\Delta_{1}$ is a $d_{1}\times d_{1}$ Gaussian matrix under $\overline{\mathbb{E}}$, the Sudakov-Fernique inequality (Lemma 5) yields
$\sqrt{n}\|\Theta_{1}\|_{2}^{-1/2}\overline{\mathbb{E}}\|\Delta_{1}\|_{2}\leq\sqrt{2}\mathbb{E}\sup_{\|u\|_{2}=\|v\|_{2}=1}(u^{\top}\xi+v^{\top}\zeta)$
$=$ $\sqrt{2}\mathbb{E}\left(\|\xi\|_{2}+\|\zeta\|_{2}\right)\leq 2\sqrt{2d_{1}}.$
It follows that
Elementary calculation shows that $\|\Delta_{1}\|_{2}$ is a $\sqrt{n}\|\Theta_{1}\|_{2}^{1/2}$-Lipschitz function. Then, by Gaussian concentration inequality for Lipschitz functions (Lemma 6),
With $x\asymp\sqrt{d_{1}}$, with probability at least $1-e^{-d_{1}}$,
(ii) Iterative refinement.
[…]
$\widehat{A}_{1}^{(m+1)}$ $=\text{LSVD}_{r_{1}}(\text{TIPUP}_{1}(\widehat{A}_{2}^{(m)})),$
$\widehat{P}_{1}^{(m+1)}$ $=\widehat{A}_{1}^{(m+1)}\widehat{A}_{1}^{(m+1)^{\top}}.$
The “noiseless” version of this update is given by
giving error-free “estimates”,
$\widehat{A}_{1}^{(m+1)}=\text{LSVD}_{r_{1}}(\Theta_{1}(\widehat{A}_{2}^{(m)})),$
where $\Theta_{1}(\widehat{A}_{2}^{(m)})$ is of rank $r_{1}$. Thus, by Davis-Kahan sin $\Theta$ theorem (Lemma 2)
To bound the numerator on the right-hand side of (S.17), we write
where for any $M_{2}\in\mathbb{R}^{d_{2}\times d_{2}}$,
$\|\text{TIPUP}_{1}(\widehat{A}_{2}^{(m)})-\Theta_{1}(\widehat{A}_{2}^{(m)})\|_{2}$
$\leq$ $\|\text{TIPUP}_{1}(A_{2})-\Theta_{1}(A_{2})\|_{2}+L^{(m)}(2K-2)\sum_{i=1}^{3}\|\Delta_{i}(M_{2})\|_{\rm S},$
where
We will also prove by induction, the denominator in (S.17) satisfies
Define the ideal version of the ratio in (S.17) as
$L_{k}^{(\text{ideal})}$ $=\frac{\|\text{TIPUP}_{1}(\odot_{j\neq k}A_{j})-\Theta_{k}(\odot_{j\neq k}A_{j})\|_{2}}{\sigma_{r_{1}}[\Theta_{k}(\odot_{j\neq k}A_{j})]},$
$L^{(\text{ideal})}$ $=\max_{1\leq k\leq K}L_{k}^{(\text{ideal})}.$
Putting together (S.15), (S.16), (S.17), (S.18), (S.19) and (S.20), we have
$L_{k}^{(m+1)}\leq L_{k}^{(\text{ideal})}+L^{(m)}\rho,$
with $\rho<1$. By induction,
$L^{(m+1)}\leq(1+\rho+\dots+\rho^{m})L^{(m)}+\rho^{m+1}L^{(0)}.$
We achieve the desired results.
Step 1.
We prove (S.18) for the $\|\Delta_{1}(M_{2})\|_{\rm S}$.
By Lemma 3 (ii), there exists $\widetilde{M}^{(\ell,\ell^{\prime})}\in\mathbb{R}^{d_{2}\times d_{2}}$ of the forms $W_{\ell}Q_{\ell^{\prime}}^{\top}$ with $W_{\ell}\in\mathbb{R}^{d_{2}\times r_{2}}$, $Q_{\ell^{\prime}}\in\mathbb{R}^{d_{2}\times r_{2}}$, $1\leq\ell,\ell^{\prime}\leq N_{d_{2}r_{2},1/8}=17^{d_{2}r_{2}}$, such that $\|\widetilde{M}^{(\ell,\ell^{\prime})}\|_{2}\leq 1$, $\text{rank}(\widetilde{M}^{(\ell,\ell^{\prime})})\leq r_{2}$ and
Elementary calculation shows that $\|\Delta_{1}(\widetilde{M}^{(\ell,\ell^{\prime})})\|_{2}$ is a $\sqrt{n}\|\Theta_{1}\|_{2}$-Lipschitz function of $\mathcal{E}_{1},\ldots,\mathcal{E}_{n}$. Employing similar argument as the initialization (Sudakov-Fernique inequality, Lemma 5), we have
Then, by Gaussian concentration inequality for Lipschitz functions (Lemma 6),
¯$\overline{\mathbb{P}}\left(\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\widetilde{M}^{(\ell,\ell^{\prime})}\mathcal{E}_{t}^{\top}\right\|_{2}-\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}\right.$
$\qquad\geq\left.\frac{1}{\sqrt{n}}\|\Theta_{1}\|_{2}^{1/2}x\right)$
$\leq$ $e^{-x^{2}/2}.$
Hence,
¯$\overline{\mathbb{P}}\left(\|\Delta_{1}(M_{2})\|_{\rm S}\geq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}+\frac{1}{\sqrt{n}}\|\Theta_{1}\|_{2}^{1/2}x\right)$
$\leq$ $N_{d_{2}r_{2},1/8}^{2}e^{-x^{2}/2}.$
Step 2.
Step 3.
Here we prove (S.19) for $\|\Delta_{3}(M_{2})\|_{\rm S}$. By Lemma 3(iii), we can find $U_{2}^{(\ell)}\in\mathbb{R}^{d_{2}\times r_{2}}$, $1\leq\ell\leq N_{d_{2}r_{2},1/8}$ such that $\|U_{2}^{(\ell)}\|_{2}\leq 1$ and
By Lemma 7,
$\mathbb{P}\left(\left\|\sum_{t=1}^{n}\mathcal{E}_{t}U_{2}^{(\ell)}U_{2}^{(\ell)\top}\mathcal{E}_{t}^{\top}-n\mathbb{E}\mathcal{E}_{t}U_{2}^{(\ell)}U_{2}^{(\ell)\top}\mathcal{E}_{t}^{\top}\right\|_{2}\right.$
$\qquad\geq\left.d_{1}+2\sqrt{d_{1}r_{2}n}+2x(\sqrt{d_{1}}+\sqrt{r_{2}n})+x^{2}\right)$
$\leq$ $2e^{-x^{2}/2}.$
$\|\Delta_{3}(M_{2})\|_{\rm S}$ $\lesssim\frac{d_{1}}{n}+\sqrt{\frac{d_{1}r_{2}}{n}}+\frac{d_{2}r_{2}}{n}+\sqrt{\frac{d_{2}r_{2}^{2}}{n^{2}}}$
$\lesssim\frac{d_{1}+d_{2}}{n}+\sqrt{\frac{d_{1}+d_{2}}{n}}.$
Step 4.
Next, we prove (S.20). Note that
$\|\Theta_{1}(\widehat{A}_{2}^{(m)})-\Theta_{1}(A_{2})\|_{2}$
$\leq\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}(\widehat{A}_{2}^{(m)}\widehat{A}_{2}^{(m)\top}-A_{2}A_{2}^{\top})A_{2}F_{t}^{\top}A_{1}^{\top}\right\|_{2}$
$\leq L^{(m)}\|\Theta_{1}\|_{2}.$
Thus, by Weyl’s inequality (Lemma 1),
when $\min_{1\leq k\leq K}\frac{\sigma_{r_{k}}[\Theta_{k}(A_{2})]}{\|\Theta_{k}(A_{2})\|_{2}}\geq 2L^{(m)}$. We prove this condition by induction. ∎
Appendix B Appendix B: Supplemental Numerical Results
Supplement to FAO Data Analysis
Studies on FAO data can reveal temporal causal links that help inform decisions, policies, and investments related to food and agriculture. The dataset has also been examined in other studies (e.g., Chepeliev (2020); Lee and Wang (2023); Chevuru et al. (2023)) to address similar issues discussed in this article. Predicting yield and production in South America based on data from East Asia, North America and Europe can aid in analyzing global climate effects. For example, El Ni no causes droughts in Asia and heavy rains in South America. Studying El Ni no’s impact on crop yields in East Asia and North America helps anticipate its effects on South American agriculture (Anderson et al. 2017). Such predictions also enhance our understanding of global market dynamics. South America and the US are top soybean producers, and changes in production on one continent influence prices and planting decisions on the other. A US drought reducing soybean yields could drive up prices, prompting South American farmers to increase cultivation, affecting future yields (Robinson et al. 2015).
  • (i)
  • (ii)
Figures 2, S.1 and S.2, visualize the comparisons between ground-truth values and predicted metrics in the two prediction tasks for FAO data. The predicted and actual quantities are represented by the color intensity of each country. A closer examination reveals that our proposed FATTNN model produces predictions that more closely align with the ground truth compared to the benchmark methods (right panel). The color gradients in the FATTNN predictions more accurately reflect the actual data, indicating superior spatial prediction performance. We have also included numerical values tabularly reported in Table S.2 and Table S.3, providing a quantitative comparison that reflects the advantages of our method.
Supplement to NYC Taxi Data Analysis
Factorized temporal tensor acts like time-varying principal components, summarizing the key information of data in reduced dimensionality. Specific interpretation of core tensor varies across datasets. For example, in taxi data, the dimension reduction from (19,19) (pick-up, drop-off districts) in raw data to (4,4) in core tensor reflects clusters of geographical regions that share similar traffic patterns, and the reduction from 8 (hours) in raw data to 2 in core tensor summarizes traffic patterns into rush hour and off-peak hour.
Figure 3 (with an enlarged version shown in Figure S.4) presents a comparison between ground-truth values and predicted pick-up and drop-off volumes using various methods. In this figure, the ground truth is depicted by the blue solid line, while our FATTNN predictions are shown as a dark red line with smaller dots. The FATTNN line closely follows the ground truth, both in shape and location, outperforming the benchmark method represented by the red line. This demonstrates our model’s enhanced ability to capture temporal patterns in the New York taxi data.
Supplement to FMRI Data Analysis
Figure S.5 presents a comparison between ground-truth values and predicted outputs for selected samples in FMRI dataset. The figure shows that the FMRI image predicted by FATTNN closely resembles the actual FMRI image shown on the left. In contrast, the images generated by TCN, LSTM, and TRL without considering tensor factor structures, shows a more blur pattern and noticeable discrepancies. The Conv-TT-LSTM also produces a plot that resembles the actual FMRI image well. The figure is also consistent with the test MSE results summarized in Table 3, that Conv-TT-LSTM yields the lowest MSE in the FMRI prediction task, followed by the FATTNN. This is not surprising as Conv-TT-LSTM is specifically designed for image prediction. We also observe a noticeable difference in computation time. Though Conv-TT-LSTM has the best prediction accuracy, it is quite computationally heavy - about 50 times slower than FATTNN.
Appendix C Appendix C: An Extended Discussion
on the Generalizations of FATTNN
In this article, we focus on tensor-on-tensor time series forecasting. The integration of TIPUP and TCN, both designed specifically for handling temporal data, makes the proposed approach a powerful tool for forecasting tensor series over time. For non-temporal data, the TIPUP-TCN-based FATTNN is still applicable for tensor-on-tensor prediction tasks, but the strengths of TIPUP and TCN may not be fully utilized. This limitation can be easily rectified by adopting proper types of tensor factor models and neural networks in the implementation of FATTNN, depending on the nature of the prediction tasks and data types involved.
Our proposed FATTNN framework is not confined to temporal data but can accommodate a variety of data types, though we presented our models for the purpose of tensor-on-tensor time series forecasting. The main idea of FATTNN is to use a low-rank tensor factor model to capture the intrinsic patterns among the observed covariates, and then proceed with a neural network to model the intricate relationships between covariates and responses. With appropriate choices of factor models and neural network architectures, the proposed FATTNN can be naturally adapted to any simple (e.g., i.i.d observations) or complex tensor-type data (e.g., image, graph, network, text, video). When handling time series tensors, we adopt TIPUP factorization and TCN to fully capitalize on their temporal nature. For higher-order tensors (e.g. $K>4$ ), Tensor-Train (Oseledets 2011) and Hierarchical Tucker decompositions (Lubich et al. 2013) may be better suited. To extend beyond time series data, different types of factor models and neural network architectures should be deployed, depending on the nature of the data. For example, for graph data, Graph Neural Network is a more suitable alternative to TCN. For text data, Transformer models could substitute for the TCN in our FATTNN framework. Our proposed FATTNN is flexible in accommodating different types of tensor factor models and deep learning architectures for tensor-on-tensor prediction using diverse data types.